3. Suppose GenericFish Corp produced 5,000 cans of herring in the fourth quarter of 2018. GenericFish has two production costs: labor and intermediate goods. GenericFish pays its workers $25 per hour, and they employed 1,000 work hours. Intermediate goods were bought from FishyStuff Inc for $4 per can. FishyStuffs only input was labor, which cost $10, 000 total. In the fourth quarter, GenericFish sold 4,000 cans to consumers at a price of $15. What was GenericFishs total contribution to GDP in Q4 of 2018? What was FishyStuffs total contribution to national income (wages and profits) in Q4 of 2018? (Round to the nearest whole dollar) FishyStus$10, 000 $60,000; NİFishyStuff-$20,000 A. GDPo4.2018 $60, 000; NI B. GDPQ4:2018 C. GDPo4:2018-$69,000; NİFishyStuff-$10,000 D. GDP24:20 18-$69,000; NİFishyStuff-$20,000 4. Consider again the scenario posed in question 3. Suppose in the first quarter of 2019, Gener- icFish produced 5,800 cans. GenericFishs wage per hour went up by 35%. GenericFish employed the same number of work hours. Also, FishyStuffs labor costs were the same as in the previous quarter. GenericFish sold to consumers all cans produced in the first quarter of 2019 and the cans left in inventory from the previous quarter (fourth quarter 2018). However, GenericFish raised the price of their canned fish to $17 after a fishy mar- keting campaign that implied anti-aging benefits associated with the consumption of their product. What were the total wages and profits for Q1 2019? A. ProfitTotal 62,850; WagesTtul=43,750 ProfitTotal 62,850; WagesTotal 39,000 C. ProfitTotal 54,850; WagesTotal 43,750 ProfitTotal 54,850; WagesTotal 39,000
Answer #1

Answer :Option C:$69000, $10000

Generic fish total contribution to GDP =Value of sold good + value of unsold good=15*4000+{(25*1000/5000)*1000+4*1000}=$69000

Fishy stuff contribution to national income =price of good sold - cost of labor =(4*5000-10000)=$10000